\begin{thebibliography}{17}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Almazrouei et~al.(2023)Almazrouei, Alobeidli, Alshamsi, Cappelli,
  Cojocaru, Debbah, Goffinet, Heslow, Launay, Malartic, Noune, Pannier, and
  Penedo]{falcon40b}
Ebtesam Almazrouei, Hamza Alobeidli, Abdulaziz Alshamsi, Alessandro Cappelli,
  Ruxandra Cojocaru, Merouane Debbah, Etienne Goffinet, Daniel Heslow, Julien
  Launay, Quentin Malartic, Badreddine Noune, Baptiste Pannier, and Guilherme
  Penedo.
\newblock {Falcon-40B}: an open large language model with state-of-the-art
  performance.
\newblock 2023.

\bibitem[Bai et~al.(2023)Bai, Bai, Chu, Cui, Dang, Deng, Fan, Ge, Han, Huang,
  Hui, Ji, Li, Lin, Lin, Liu, Liu, Lu, Lu, Ma, Men, Ren, Ren, Tan, Tan, Tu,
  Wang, Wang, Wang, Wu, Xu, Xu, Yang, Yang, Yang, Yang, Yao, Yu, Yuan, Yuan,
  Zhang, Zhang, Zhang, Zhang, Zhou, Zhou, Zhou, and Zhu]{bai2023qwen}
Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan,
  Wenbin Ge, Yu~Han, Fei Huang, Binyuan Hui, Luo Ji, Mei Li, Junyang Lin, Runji
  Lin, Dayiheng Liu, Gao Liu, Chengqiang Lu, Keming Lu, Jianxin Ma, Rui Men,
  Xingzhang Ren, Xuancheng Ren, Chuanqi Tan, Sinan Tan, Jianhong Tu, Peng Wang,
  Shijie Wang, Wei Wang, Shengguang Wu, Benfeng Xu, Jin Xu, An~Yang, Hao Yang,
  Jian Yang, Shusheng Yang, Yang Yao, Bowen Yu, Hongyi Yuan, Zheng Yuan,
  Jianwei Zhang, Xingxuan Zhang, Yichang Zhang, Zhenru Zhang, Chang Zhou,
  Jingren Zhou, Xiaohuan Zhou, and Tianhang Zhu.
\newblock {Qwen Technical Report}.
\newblock \emph{arXiv preprint}, arXiv:2309.16609, 2023.

\bibitem[Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan,
  Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess,
  Clark, Berner, McCandlish, Radford, Sutskever, and
  Amodei]{NEURIPS2020_1457c0d6}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla
  Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell,
  Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon
  Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris
  Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess,
  Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever,
  and Dario Amodei.
\newblock {Language Models are Few-Shot Learners}.
\newblock In \emph{Proceedings of NeurIPS}, 2020.

\bibitem[Dao et~al.(2022)Dao, Fu, Ermon, Rudra, and
  R{\'e}]{dao2022flashattention}
Tri Dao, Daniel~Y. Fu, Stefano Ermon, Atri Rudra, and Christopher R{\'e}.
\newblock {Flash{A}ttention: Fast and Memory-Efficient Exact Attention with
  {IO}-Awareness}.
\newblock In \emph{Proceedings of NeurIPS}, 2022.

\bibitem[Ding et~al.(2023)Ding, Chen, Xu, Qin, Zheng, Hu, Liu, Sun, and
  Zhou]{ding2023enhancing}
Ning Ding, Yulin Chen, Bokai Xu, Yujia Qin, Zhi Zheng, Shengding Hu, Zhiyuan
  Liu, Maosong Sun, and Bowen Zhou.
\newblock Enhancing chat language models by scaling high-quality instructional
  conversations, 2023.

\bibitem[Gemma-Team et~al.(2024)Gemma-Team, Mesnard, Hardin, Dadashi,
  Bhupatiraju, Pathak, Sifre, Rivière, Kale, Love, Tafti, Hussenot, Chowdhery,
  Roberts, Barua, Botev, Castro-Ros, Slone, Héliou, Tacchetti, Bulanova,
  Paterson, Tsai, Shahriari, Lan, Choquette-Choo, Crepy, Cer, Ippolito, Reid,
  Buchatskaya, Ni, Noland, Yan, Tucker, Muraru, Rozhdestvenskiy, Michalewski,
  Tenney, Grishchenko, Austin, Keeling, Labanowski, Lespiau, Stanway, Brennan,
  Chen, Ferret, Chiu, Mao-Jones, Lee, Yu, Millican, Sjoesund, Lee, Dixon, Reid,
  Mikuła, Wirth, Sharman, Chinaev, Thain, Bachem, Chang, Wahltinez, Bailey,
  Michel, Yotov, Sessa, Chaabouni, Comanescu, Jana, Anil, McIlroy, Liu,
  Mullins, Smith, Borgeaud, Girgin, Douglas, Pandya, Shakeri, De, Klimenko,
  Hennigan, Feinberg, Stokowiec, hui Chen, Ahmed, Gong, Warkentin, Peran,
  Giang, Farabet, Vinyals, Dean, Kavukcuoglu, Hassabis, Ghahramani, Eck,
  Barral, Pereira, Collins, Joulin, Fiedel, Senter, Andreev, and
  Kenealy]{gemmateam2024gemma}
Gemma-Team, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju,
  Shreya Pathak, Laurent Sifre, Morgane Rivière, Mihir~Sanjay Kale, Juliette
  Love, Pouya Tafti, Léonard Hussenot, Aakanksha Chowdhery, Adam Roberts,
  Aditya Barua, Alex Botev, Alex Castro-Ros, Ambrose Slone, Amélie Héliou,
  Andrea Tacchetti, Anna Bulanova, Antonia Paterson, Beth Tsai, Bobak
  Shahriari, Charline~Le Lan, Christopher~A. Choquette-Choo, Clément Crepy,
  Daniel Cer, Daphne Ippolito, David Reid, Elena Buchatskaya, Eric Ni, Eric
  Noland, Geng Yan, George Tucker, George-Christian Muraru, Grigory
  Rozhdestvenskiy, Henryk Michalewski, Ian Tenney, Ivan Grishchenko, Jacob
  Austin, James Keeling, Jane Labanowski, Jean-Baptiste Lespiau, Jeff Stanway,
  Jenny Brennan, Jeremy Chen, Johan Ferret, Justin Chiu, Justin Mao-Jones,
  Katherine Lee, Kathy Yu, Katie Millican, Lars~Lowe Sjoesund, Lisa Lee, Lucas
  Dixon, Machel Reid, Maciej Mikuła, Mateo Wirth, Michael Sharman, Nikolai
  Chinaev, Nithum Thain, Olivier Bachem, Oscar Chang, Oscar Wahltinez, Paige
  Bailey, Paul Michel, Petko Yotov, Pier~Giuseppe Sessa, Rahma Chaabouni,
  Ramona Comanescu, Reena Jana, Rohan Anil, Ross McIlroy, Ruibo Liu, Ryan
  Mullins, Samuel~L Smith, Sebastian Borgeaud, Sertan Girgin, Sholto Douglas,
  Shree Pandya, Siamak Shakeri, Soham De, Ted Klimenko, Tom Hennigan, Vlad
  Feinberg, Wojciech Stokowiec, Yu~hui Chen, Zafarali Ahmed, Zhitao Gong, Tris
  Warkentin, Ludovic Peran, Minh Giang, Clément Farabet, Oriol Vinyals, Jeff
  Dean, Koray Kavukcuoglu, Demis Hassabis, Zoubin Ghahramani, Douglas Eck,
  Joelle Barral, Fernando Pereira, Eli Collins, Armand Joulin, Noah Fiedel,
  Evan Senter, Alek Andreev, and Kathleen Kenealy.
\newblock {Gemma: Open Models Based on Gemini Research and Technology}.
\newblock \emph{arXiv preprint}, arXiv:2403.08295, 2024.

\bibitem[Jiang et~al.(2023)Jiang, Sablayrolles, Mensch, Bamford, Chaplot,
  de~las Casas, Bressand, Lengyel, Lample, Saulnier, Lavaud, Lachaux, Stock,
  Scao, Lavril, Wang, Lacroix, and Sayed]{jiang2023mistral}
Albert~Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford,
  Devendra~Singh Chaplot, Diego de~las Casas, Florian Bressand, Gianna Lengyel,
  Guillaume Lample, Lucile Saulnier, Lélio~Renard Lavaud, Marie-Anne Lachaux,
  Pierre Stock, Teven~Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix,
  and William~El Sayed.
\newblock {Mistral 7B}.
\newblock \emph{arXiv preprint}, arXiv:2310.06825, 2023.

\bibitem[Kwon et~al.(2023)Kwon, Li, Zhuang, Sheng, Zheng, Yu, Gonzalez, Zhang,
  and Stoica]{kwon2023efficient}
Woosuk Kwon, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody~Hao Yu,
  Joseph~E. Gonzalez, Hao Zhang, and Ion Stoica.
\newblock {Efficient Memory Management for Large Language Model Serving with
  PagedAttention}.
\newblock In \emph{Proceedings of SIGOPS}, 2023.

\bibitem[Li et~al.(2023)Li, Koto, Wu, Aji, and Baldwin]{li2023bactrianx}
Haonan Li, Fajri Koto, Minghao Wu, Alham~Fikri Aji, and Timothy Baldwin.
\newblock {Bactrian-X : A Multilingual Replicable Instruction-Following Model
  with Low-Rank Adaptation}.
\newblock \emph{arXiv preprint}, arXiv:2305.15011, 2023.

\bibitem[Lin \& Chen(2023)Lin and Chen]{lin-chen-2023-llm}
Yen-Ting Lin and Yun-Nung Chen.
\newblock {{LLM}-Eval: Unified Multi-Dimensional Automatic Evaluation for
  Open-Domain Conversations with Large Language Models}.
\newblock In Yun-Nung Chen and Abhinav Rastogi (eds.), \emph{Proceedings of
  NLP4ConvAI}, pp.\  47--58, 2023.

\bibitem[Nguyen et~al.(2023)Nguyen, Tran, Nguyen, and
  Vu]{nguyen-etal-2023-vigptqa}
Minh~Thuan Nguyen, Khanh~Tung Tran, Nhu~Van Nguyen, and Xuan-Son Vu.
\newblock {V}i{GPTQA} - state-of-the-art {LLM}s for {V}ietnamese question
  answering: System overview, core models training, and evaluations.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in
  Natural Language Processing: Industry Track}, pp.\  754--764, 2023.

\bibitem[Press et~al.(2022)Press, Smith, and Lewis]{press2022train}
Ofir Press, Noah Smith, and Mike Lewis.
\newblock {Train Short, Test Long: Attention with Linear Biases Enables Input
  Length Extrapolation}.
\newblock In \emph{Proceedings of ICLR}, 2022.

\bibitem[Team(2023)]{MosaicML2023Introducing}
MosaicML~NLP Team.
\newblock Introducing mpt-7b: A new standard for open-source, commercially
  usable llms, 2023.
\newblock URL \url{www.mosaicml.com/blog/mpt-7b}.
\newblock Accessed: 2023-05-05.

\bibitem[Touvron et~al.(2023{\natexlab{a}})Touvron, Lavril, Izacard, Martinet,
  Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, Rodriguez, Joulin, Grave,
  and Lample]{touvron2023llama}
Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne
  Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro,
  Faisal Azhar, Aurelien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume
  Lample.
\newblock {LLaMA: Open and Efficient Foundation Language Models}.
\newblock \emph{arXiv preprint}, arXiv:2302.13971, 2023{\natexlab{a}}.

\bibitem[Touvron et~al.(2023{\natexlab{b}})Touvron, Martin, Stone, Albert,
  Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, Bikel, Blecher,
  Ferrer, Chen, Cucurull, Esiobu, Fernandes, Fu, Fu, Fuller, Gao, Goswami,
  Goyal, Hartshorn, Hosseini, Hou, Inan, Kardas, Kerkez, Khabsa, Kloumann,
  Korenev, Koura, Lachaux, Lavril, Lee, Liskovich, Lu, Mao, Martinet, Mihaylov,
  Mishra, Molybog, Nie, Poulton, Reizenstein, Rungta, Saladi, Schelten, Silva,
  Smith, Subramanian, Tan, Tang, Taylor, Williams, Kuan, Xu, Yan, Zarov, Zhang,
  Fan, Kambadur, Narang, Rodriguez, Stojnic, Edunov, and
  Scialom]{touvron2023llama2}
Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine
  Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale,
  Dan Bikel, Lukas Blecher, Cristian~Canton Ferrer, Moya Chen, Guillem
  Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller,
  Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar
  Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa,
  Isabel Kloumann, Artem Korenev, Punit~Singh Koura, Marie-Anne Lachaux,
  Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier
  Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew
  Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan
  Silva, Eric~Michael Smith, Ranjan Subramanian, Xiaoqing~Ellen Tan, Binh Tang,
  Ross Taylor, Adina Williams, Jian~Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan
  Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien
  Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom.
\newblock {Llama 2: Open Foundation and Fine-Tuned Chat Models}.
\newblock \emph{arXiv preprint}, arXiv:2307.09288, 2023{\natexlab{b}}.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez,
  Kaiser, and Polosukhin]{NIPS2017_7181}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin.
\newblock {Attention is All you Need}.
\newblock In \emph{Proceedings of NIPS}, pp.\  5998--6008, 2017.

\bibitem[Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, Davison, Shleifer, von Platen, Ma, Jernite, Plu, Xu,
  Scao, Gugger, Drame, Lhoest, and Rush]{wolf-etal-2020-transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe
  Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien
  Plu, Canwen Xu, Teven~Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest,
  and Alexander~M. Rush.
\newblock {Transformers: State-of-the-Art Natural Language Processing}.
\newblock In \emph{Proceedings of EMNLP: System Demonstrations}, pp.\  38--45,
  2020.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2023,times}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage[hidelinks]{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage{graphicx}
\usepackage{multirow}
\usepackage{tabularx}
\usepackage{arydshln}
\usepackage[T5]{fontenc}
\iclrfinaltrue

\title{PhoGPT: \\ Generative Pre-training for Vietnamese}

\author{Dat Quoc Nguyen, Linh The Nguyen, Chi Tran, Dung Ngoc Nguyen, Dinh Phung, Hung Bui  \\
VinAI Research, Hanoi, Vietnam\\
  \small{\texttt{\{v.datnq9, v.linhnt140, v.chitb, v.dungnn28, v.dinhpq2, v.hungbh1\}@vinai.io}}
}

\begin{document}

\maketitle

\begin{abstract}
We open-source a state-of-the-art 4B-parameter generative model series for Vietnamese, which includes the base pre-trained monolingual model PhoGPT-4B and its chat variant, PhoGPT-4B-Chat. The base model, PhoGPT-4B, with exactly 3.7B parameters, is pre-trained from scratch on a Vietnamese corpus of 102B tokens, with an 8192 context length, employing a vocabulary of 20480 token types. The chat variant, PhoGPT-4B-Chat, is the modeling output obtained by fine-tuning PhoGPT-4B on a dataset of 70K instructional prompts and their responses, along with an additional 290K conversations. In addition, we also demonstrate its superior performance compared to previous open-source models. 
Our PhoGPT models are available at: \url{https://github.com/VinAIResearch/PhoGPT}
\end{abstract}

\section{Introduction}

Undoubtedly, the success of large language models (LLMs), particularly decoder-only transformer-based generative models such as ChatGPT/GPT-4, 
LLaMA/LLaMA-2 \citep{touvron2023llama,touvron2023llama2}, Mistral \citep{jiang2023mistral} and Falcon \citep{falcon40b},  stands out as one of the most significant achievements in recent AI research and development. However, that success has largely been limited to English. 

For Vietnamese, we now release the 4B-parameter base pre-trained monolingual model, PhoGPT-4B, along with its chat variant, PhoGPT-4B-Chat, as open-source. We pre-train the base model PhoGPT-4B from scratch on a Vietnamese corpus of 102B tokens for two epochs, with an 8192 context length. Here, it employs a Vietnamese-specific byte-level BPE tokenizer with a vocabulary of 20480 tokens. We further fine-tune the base model on a dataset of 70K instructional prompts and their responses, along with an additional 290K conversations, resulting in the chat variant, PhoGPT-4B-Chat. We demonstrate its strong performance compared to previous closed-source and open-source 7B-parameter models. 

Our goal is to provide comprehensive and powerful LLMs for Vietnamese, facilitating future research and applications in generative Vietnamese NLP. Our PhoGPT can be used with popular libraries such as ``transformers'' \citep{wolf-etal-2020-transformers}, ``vllm'' \citep{kwon2023efficient} and ``llama.cpp''.\footnote{\url{https://github.com/ggerganov/llama.cpp}}




\section{PhoGPT}

\subsection{PhoGPT-4B: Model architecture and Pre-training}

PhoGPT-4B is a Transformer decoder-based model \citep{NEURIPS2020_1457c0d6,NIPS2017_7181},  which incorporates (Triton) flash attention \citep{dao2022flashattention} and ALiBi \citep{press2022train} for context length extrapolation. We train a Vietnamese-specific byte-level BPE tokenizer with a vocabulary of 20480 tokens using the ``tokenizers'' library.\footnote{\url{https://github.com/huggingface/tokenizers}}  In addition, we use a ``max\_seq\_len'' of 8192, ``d\_model'' of 3072, ``n\_heads'' of 24 and ``n\_layers'' of 32, resulting in a model size of 3.7B  parameters ($\sim$4B).  Utilizing the Mosaicml ``llm-foundry'' library \citep{MosaicML2023Introducing},\footnote{\url{https://github.com/mosaicml/llm-foundry}} 
we pre-train PhoGPT-4B from scratch  on a 482GB deduplicated and cleaned pre-training corpus of Vietnamese texts ($\sim$102B tokens) for {two epochs}. Our pre-training Vietnamese corpus consists of:\footnote{Last crawling/cutoff date: 31/05/2023.}

\begin{itemize}
\item 1GB of Wikipedia texts (version 20/05/2023);
\item 1.5GB of medical-related texts crawled from a wide range of publicly available and medical domain-specific websites such as medical journals and universities;\footnote{The content extracted from these sources contains no private data about the patients.}
\item 3GB of publicly available books spanning a range of genres;
\item 12GB of legal data crawled from \url{thuvienphapluat.vn}
 and \url{lawnet.vn};
\item a 40GB variant of the "binhvq" news corpus (version 21/05/2021);\footnote{\url{https://github.com/binhvq/news-corpus}}
\item an 88GB variant of the Vietnamese OSCAR-2301 subset;\footnote{\url{https://huggingface.co/datasets/oscar-corpus/OSCAR-2301}}
\item a 336GB variant of the Vietnamese mC4 subset.\footnote{\url{https://huggingface.co/datasets/allenai/c4/tree/mC4_3.1.0}}
\end{itemize}

\subsection{PhoGPT-4B-Chat: Supervised fine-tuning}

We then fine-tune the base pre-trained PhoGPT-4B using a dataset consisting of 70K instructional prompts and their responses, along with an additional 290K conversations, constructed by concatenating the following sources:

\begin{itemize}
\item 500 instructional prompt and response pairs for poem writing, 500 for essay writing, 500 for spelling correction,  500 for single-document summarization and 1000  for context-based question answering;
\item 67K instructional prompt and response pairs from the Vietnamese subset of Bactrian-X \citep{li2023bactrianx};
\item 20K Vietnamese-translated ChatAlpaca  conversations;\footnote{\url{https://github.com/cascip/ChatAlpaca}}
\item 40K Vietnamese-translated ShareGPT conversations (without code and mathematics);\footnote{\url{https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered}} 
\item 230K Vietnamese-translated UltraChat conversations \citep{ding2023enhancing};\footnote{\url{https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k} (including both train \& test sets)}
\end{itemize}
 
The resulting fine-tuned model is named PhoGPT-4B-Chat.

\section{Evaluation}


We compare PhoGPT-4B-Chat with the closed-source models GPT-4-0125-preview, GPT-3.5-turbo and Gemini Pro 1.0, as well as other open-source models, including:

\begin{itemize}
\item Vistral-7B-Chat is the modeling output obtained by continually pre-training Mistral-7B \citep{jiang2023mistral} on a diverse corpus of Vietnamese texts and then performing supervised fine-tuning using a diverse instructional and conversational dataset.\footnote{\url{https://huggingface.co/Viet-Mistral/Vistral-7B-Chat}}

\item SeaLLM-7B-v2 is the modeling output obtained by continually pre-training Mistral-7B on a multilingual corpus from Southeast Asian (SEA) languages, including Vietnamese, and then performing supervised fine-tuning using instructional question and answer pairs.\footnote{\url{https://huggingface.co/SeaLLMs/SeaLLM-7B-v2}}

\item Sailor-7B-Chat\footnote{\url{https://huggingface.co/sail/Sailor-7B-Chat}} and Sailor-4B-Chat \footnote{\url{https://huggingface.co/sail/Sailor-4B-Chat}} are the modeling outputs obtained by continually pre-training Qwen1.5-7B and Qwen1.5-4B \citep{bai2023qwen} on 400B tokens from SEA languages, including Vietnamese, and then performing supervised fine-tuning using instructional question and answer pairs.

\item VBD-LLaMA2-7B-50b-Chat is the modeling output obtained by continually pre-training LLaMA-2-7B \citep{touvron2023llama2} on a data combination of 40B Vietnamese tokens and 16B English tokens and then performing supervised fine-tuning using 2M instructional and conversational samples.\footnote{\url{https://huggingface.co/LR-AI-Labs/vbd-llama2-7B-50b-chat}}

\item VinaLLaMA-7B-Chat is the modeling output obtained by continually pre-training LLaMA-2-7B on a data combination of 100B English tokens, 230B Vietnamese tokens (from books and news), and 500B automatically generated Vietnamese tokens, and then performing supervised fine-tuning using 1M instructional and conversational samples.\footnote{\url{https://huggingface.co/vilm/vinallama-7b-chat}}

\item Gemma-7B-it is the 7B instruct version of the Gemma model \citep{gemmateam2024gemma}.\footnote{\url{https://huggingface.co/google/gemma-7b-it}}

\end{itemize}

\begin{table*}[!t]
\centering
\begin{tabular}{l|c|c}
\hline 
\textbf{Model} & \textbf{All truthful questions} & \textbf{Vietnam-specific} \\
\hline
PhoGPT-4B-Chat	& \underline{41.7	(83 / 199)}	& \textbf{43.5	(64 / 147)} \\
GPT-4-0125-preview		& \textbf{44.7	(89 / 199)}		& 39.5	(58 / 147)  \\
GPT-3.5-turbo		& 29.1	(58 / 199)		& 22.4	(33 / 147)  \\
Gemini Pro 1.0		& 39.7	(79 / 199)		& 34.7	(51 / 147)  \\
Vistral-7B-Chat		& 41.2	(82 / 199)		& \underline{42.9	(63 / 147)}  \\
Sailor-7B-Chat		& 28.6	(57 / 199)		& 27.9	(41 / 147)  \\
Sailor-4B-Chat		& 15.6	(31 / 199)		& 14.3	(21 / 147)  \\
SeaLLM-7B-v2		& 20.6	(41 / 199)		& 13.6	(20 / 147)  \\
VBD-Llama2-7B-50B-Chat		& 15.6	(31 / 199)		& 10.9	(16 / 147)  \\
Vinallama-7B-Chat		& 11.1	(22 / 199)		& 8.2	(12 / 147)  \\
Gemma-7B-it		& 8.0	(16 / 199)		& 6.1	(9 / 147)  \\
\hline 
\end{tabular}
\caption{Obtained results.}
\label{table:results}
\end{table*}


Our empirical study employs the Vietnamese truthful question-answering dataset ViTruthfulQA \citep{nguyen-etal-2023-vigptqa}, comprising 199 questions.\footnote{The original dataset consists of 213 questions; however, 14 of them are not questions or are unclear, e.g. ``Tên gọi nào không phải là một loại trái cây Việt Nam?'' (Which name is not a Vietnamese fruit?). Therefore, we remove them, resulting in a final evaluation set of 199 questions.} Each of the 199 questions is fed into 11 experimental models to generate responses, which are then anonymously shuffled. Here, we utilize the greedy search decoding method, which is more suitable for LLM comparison \citep{lin-chen-2023-llm}. Two annotators independently assess each generated response on whether the response is correct or not. A response is annotated as ``correct'' only if it contains the accurate answer for the corresponding question without any hallucinated information. We host a discussion session with the annotators to resolve annotation conflicts. 


Table \ref{table:results} presents the accuracy results obtained, showing that overall, our 4B-parameter PhoGPT-4B-Chat is highly competitive compared to the closed-source GPT-4 model, yielding better accuracy scores than the remaining closed-source models GPT-3.5-turbo and Gemini Pro 1.0, as well as all open-source baselines. 
Furthermore, when it comes to 147 out of the 199 questions that specifically ask for information related to Vietnam, PhoGPT-4B-Chat achieves the highest accuracy. 

\section{Conclusion}

We have introduced state-of-the-art open-source  4B-parameter LLMs for Vietnamese, including the base pre-trained PhoGPT-4B and its chat variant, PhoGPT-4B-Chat. 
We hope that these models will foster future research and applications of Vietnamese LLMs.

\section*{Limitations}

PhoGPT has certain limitations. For example, it is not good at tasks involving reasoning, coding or mathematics. PhoGPT may generate harmful, hate speech, biased responses, or answer unsafe questions. Users should be cautious when interacting with PhoGPT that can produce factually incorrect output.

\section*{Acknowledgments}

We extend our thanks to Nhung Nguyen (v.nhungnt89@vinai.io) for crawling and pre-processing health data and to Thien Huu Nguyen (v.thiennh4@vinai.io) for the initial discussions.

\bibliographystyle{iclr2023_conference}
\bibliography{refs}

\end{document}